OpenStudy (anonymous):

5 + 4 • (8 – 6)2 1. (8-6) = 2 2. 2 • 2= 4 3. 5+4•4 16 4. 5+16= 21 not sure if i got it right?

OpenStudy (theviper):

Which one is it 1 or 2?? 1. 5 + 4 • (8 – 6)2 \[2.\space5 + 4 • (8 – 6)^2\]

OpenStudy (theviper):

If it is first one then u r right :)

OpenStudy (anonymous):

If both of them is your question then also you are right...

OpenStudy (anonymous):

Because : \(2 \times 2 = 4\) and \(2^2 = 4\) both result into 4..

OpenStudy (anonymous):

And particularly I think your question is this : \[5 + 4 \cdot (8-6)^2\]

OpenStudy (anonymous):

And particularly I also think that your answer is right.. \(\huge \color{blue}\checkmark\)